OpenStudy (anonymous):

What is the slope of the line (-3,-2) and (3,-1)?

OpenStudy (michele_laino):

the requested slope m is given by the subsequent formula: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}}\]

OpenStudy (anonymous):

I'm really bad with negatives and positives need major help I know the damn formula tho.

OpenStudy (michele_laino):

where (x1, y1)=(-3,-2) and (x2,y2)=3,-1)

OpenStudy (michele_laino):

hint: after a substitution, we get: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 1 - \left( { - 2} \right)}}{{3 - \left( { - 3} \right)}} = ...?\]

OpenStudy (anonymous):

-1/-6= 1/6?!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Thank you <3

OpenStudy (anonymous):

For (-2,3) and (1,-3) would it be 0/-3=0 ?!

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Yup and I got 0/-3 which is 0

OpenStudy (michele_laino):

hint: we have: x1=-2, y1=3 x2=1, y2=-3 so we can write: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{ - 3 - 3}}{{1 - \left( { - 2} \right)}} = ...?\]

OpenStudy (anonymous):

Ughhh it wasn't right

OpenStudy (anonymous):

-2 ?!

OpenStudy (michele_laino):

we know the exact result at least: m=-2 that's right!!

OpenStudy (anonymous):

Thank you 😄😄

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

(-3,2) and (1,2) I got 0/3=0 :/

OpenStudy (anonymous):

@Michele_Laino